Solve the systems of equations by the substitution method:
y=9x+7
3x-2y=1
Thanks,
Vanessa

Answer by ankor@dixie-net.com(22740)

Solve the systems of equations by the substitution method:
y = 9x + 7
3x - 2y = 1
:
This isn't hard, we know y = (9x+7), replace y with this in the 2nd equation:
3x - 2(9x+7) = 1
3x - 18x - 14 = 1
-15x = 1 + 14
-15x = 15
x = -1
;
I'll let you find y, use the 1st equation, replace x with -1
:
Check your solutions in the 2nd equation